Sammamish Brick, Inc., manufactures bricks using clay depositd on the company's property. Raw clays are blended and then extruded into molds to form unfired bricks. The unfired bricks are then stacked onto movable metal platforms and rolled into the kiln where they are fired until dry. The dried bricks are then packaged and shipped to retail outlets and contractors. The bottleneck in the production process is the kiln, which is available for 2,850 hours per year. Data concerning the company's four main products appear below. Products are sold by the pallet.

4. The kiln could be operated for more than 2,850 hours per year by running it after normal working hours. Up to how much per hour should the company be willing to pay in overtime wages, energy costs, and other incremental costs to operate the kiln additional hours?

per hour Maximum amount to operate the kiln

5. The company is considering introducing a new product, glazed venetian bricks, whose variable cost would be $865 per pallet and that would require 10 hours in the kiln per pallet. What is the minimum acceptable selling price for this new product?

Time required (hours) Selling price 0 $

6. Salespersons are currently paid a commission of 5% of gross revenues. Will this motivate the salespersons to make the right choices concerning which products to sell most aggressively? Yes No

Traditional Textured Facing $1,396 $650 130 10 Cinder Block $652 $416 120 4 Roman Brick $946 $485 130 5 Brick $817 $510 110 10 Gross revenue per pallet Contribution margin per pallet Annual demand (pallets) Hours required in the kiln per pallet No fixed costs could be avoided by modifying how much is produced of any product. Required: 1a. Calculate the total hours required in the kiln to satisfy demand for all products. Traditional Textured Cinder Roman Total Brick Facing Block Brick Total hours required in the drying kiln 0
1b. Is there sufficient capacity in the kiln to satisfy demand for ll products? Yes O No 2a. Calculate the profitability index for the company's four products. Traditional Textured Cinder Roman Facing Block Brick Brick
2b. What is the production plan for the year that would maximize the company's profit? Cinder Block pallets pallets Roman Brick Textured Facing pallets Traditional Brick pallets
3. What would be the total contribution margin for the production plan you have proposed? Traditional Textured Cinder Roman Brick Brick Facing Block Total
per hour Maximum amount to operate the kiln
Time required (hours) Selling price 0 $
Answer #1

Answer 1 a

Total Hours in Kiln = Annual Demand* NO. of hours per pallet

Hence Total Hours for Traditional Brick = 110*10 = 1100

Total Hours for Textured Facing = 130*10 = 1300

Total Hours for Cinder Block = 120*4 = 480

Total Hours for Roman Brick = 130*5 = 650

Hence total hours = 1100+1300+480+650 = 3530 hours

Answer 1 b

Since total hours required is more than 2850 hours, hence the capacity is insufficient

Answer 2

Profitability Index= Contribution Margin/ Gross Revenue

Profitability Index for Traditional Brick = 510/817 = 0.62

Profitability Index for Textured Facing = 650/1396 = 0.47

Profitability Index for Cinder Block = 416/652 = 0.64

Profitability Index for Roman Brick = 485/946 = 0.51

Answer 2b

For Production plan, it is preferable to produce item that has a contribution margin per hour of kiln highest. Hence the profits will be maximized

Contribution per hour = Contribution margin/ No. of hours required in Kiln

Contribution per hour for Traditional Brick = 510/10 = 51

Contribution per hour for Textured Facing = 650/10 = 65

Contribution per hour for Cinder Block = 416/4 = 104

Contribution per hour for Roman Brick = 485/5 = 97

Since Cinder Block has highest contribution per hour, first cinder Block will be manufactured which will consume 480 hours

Next highest contribution per hour is for Roman Bock, hence next roman Brick will be manufactured which will consume another 650 hours

Next highest contribution per hour is for Textured Facing, hence next Textured Facing will be manufactured which will consume another 1300 hours

Total hours consumed = 480+650+1300=2430 hours

Remaining hours = 2850-2430=420 hours

Hence Traditional Bricks manufactured will be 420/10 = 42

Hence Answer is as follows:

Cinder Block Roman Brick Textured Facing Traditional Brick pallets pallets pallets pallets 120 130 130 42

Answer 3

Total Contribution Margin = Contribution Margin per unit* No of units

Contribution Margin for Traditional Brick = 510*42= 21420

Contribution Margin for Textured Facing = 650*130 = 84500

Contribution Margin for Cinder Block = 416*120 = 49920

Contribution Margin for Roman Brick = 485*130 = 63050

Total Contribution Margin = 21420+84500+49920+63050= $218890

Answer 4

Additional Hours required to meet all demand = 3530-2850 = 680 hours

In this 680 hours, 680/10 = 68 Traditional Bricks will be manufactured with Contribution MArgin of 510

Hence total additional contribution margin = 510*68= $34680

hence maximum per hour that can be paid =34680/680=$51

answer 5

Overall Contribution Margin = Total contribution margin/ NO. of hours = 218890/2850 = 76.80

Hence If new product is taking 10 hours it should have a contribution margin of 76.8*10 = $768

Hence selling price should be atleast 865+768 = $1633

Answer 6

No

Keeping commission on gross revenue basis will motivate sales person in selling item with highest gross revenue hence they will target products with highest gross revenue first. But company would be more interested to sell Cinder Block and Roman Brick as they have larger contribution margin per hour but has lower gross revenue.